# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v3

@@ 4 Conclusion
In particular, this structure appears indicative of the maximal symmetry group to which the function algebraically belongs. This is achieved through comparisons of various activation functions with differing maximal symmetries, indicating that symmetry is the primary predictor, rather than the specifics of the chosen function. This empirically motivates a thorough investigation of all primitives in terms of their resultant inductive biases when interacting with general networks, previously encouraged in Bird (2025a).
Tanh-like activation functions are chosen in two ways. The first is when their form is defined through maximally satisfying a permutation and sign-flip algebraic equivariance of the standard basis, termed the hyperoctahedral group. This function definition results in representational clusters about standard distinguished directions, which accumulate through training, progressively approximating a discrete direction. Whereas, functional forms which maximally abide by an orthogonal algebraic equivariance produce apparent smooth continua of representations. This indicates that the hyperoctahedral discrete group is inducing an approximately-discrete structure to emerge in representation space, implying the functional form is acting as a quantising bias on the representations. A similar result is found for algebraic permutation symmetry standard Leaky-ReLU, which also produces a quantising bias, whereas isotropic Leaky-ReLU tends to produce smooth continua too. This is highly indicative of the algebraic symmetry corresponding to the bias, as opposed to solely function specifics, since Tanh-like and Leaky-ReLU-like functions differ substantially in analytical qualities. This empirically validates an earlier hypothesis Bird (2025a).
@@
Consequently, the empirical results strongly support the prediction that choices in activation function form are not benign, but impart unintended inductive biases corresponding to their symmetry properties, which are sufficient in many cases to strongly influence representations, confirming a previous prediction Bird (2025a). The phenomenon is expected to be systematic for more general anisotropic primitive forms, and follow-up studies can investigate this.
@@
The empirical evidence thus far strongly suggests the link between algebraic symmetries and inductive biases on representations. If such a finding holds over further analyses and is determined to be generally valid over a broad scope of primitives and models, then one might conjecture that Algebraic Symmetries of Functions Utilised in Models Tend to Propagate Their Geometry Into Emergent Representational Structures Through Training as a potential principle for representational geometry. Despite the evidence presented in this paper, this general statement remains a tentative conjecture at the current time, requiring substantial further investigation. Refinement of this statement and detailed mathematical modelling can continue to be evaluated. For example, this may not be expected to hold where the symmetry cannot ‘act’ on the network, such as in the case of algebraic end-to-end symmetries Bronstein et al. (2021) which use it to transform representations with the data structure — unless this is reframed as a side-effect of such an inductive bias.
@@
Additionally, combinations of network primitives are expected to exert influence on the representations of larger models, potentially in an interacting, non-trivial hierarchical manner. Understanding this is a pivotal goal of Taxonomic Deep Learning’s approach Bird (2025a). This is partially demonstrated in the deeper models discussed in the appendices. Yet, overall, these potentially conflating and obscuring influences motivated the first-principles, minimalistic, and ablation-study approach undertaken in this paper, which isolates the cause to individual primitives in otherwise discretising bias-free autoencoders, as opposed to larger models.
@@
This work also corroborates the findings from the Spotlight Resonance Method study Bird (2025b) and provides a more fundamental causal framework for its observed alignment phenomenon. Coupled with the findings of this prior study, this work suggests a function-driven, symmetry-based causal mechanism for the emergence of several interpretability phenomena, suggesting that they are predicated upon the contemporary anisotropic functional forms. This representational response to anisotropic functional forms appears to explain the downstream emergence of grandmother neurons, generally discrete linear features, axis-alignment, and the representational superposition of discrete features, amongst other phenomena. Hence, axis-alignment and grandmother neurons are shown to be phenomena stimulated by design choices, rather than being spontaneous and fundamental to deep learning as a paradigm. Whilst the other phenomena appear to be strongly dependent upon such choices. Overall, this strongly supports the emergence of axis-aligned findings that are frequently observed Bau et al. (2017); Olah et al. (2019) and suggests that structural asymmetry in distributions can emerge, particularly from certain functional forms. This is considered as opposing evidence for representational uniformity Szegedy et al. (2014) or spontaneous production of axis-alignment. Together, these results may offer explanatory mechanisms to generalise to further representational geometry and interpretability phenomena.
@@
The isotropic deep learning paper Bird (2025a) also predicts that this phenomenon extends to semantics. These results demonstrate the formation of approximately discrete representations, which, when combined with the spotlight resonance method’s attribution of semantics to these representations, are considered to validate the semantic prediction as well. An end-to-end analysis could also be undertaken in future work. This direct mode for the formation of grandmother neurons is currently limited to deep learning and does not establish an empirical link to biological systems. Nevertheless, it could be used to formulate an interdisciplinary hypothesis.
In addition, the Privileged-Plane Projective method offers an alternative tool for determining alignment phenomena in representational spaces, while retaining both angular and radial information. This may be beneficial where visualisation of the distribution is desirable, but SRM may still be better suited to statistical analysis techniques.
@@
Moreover, the prediction that isotropic functional forms eliminate task-agnostic discretising structure is validated Bird (2025a). The presence of discrete-like structure is also suggested to be often circumstantially pathological Bird (2025a). Preliminary results support this statement for autoencoding networks. Isotropic formulations are shown to enable the systematic elimination of these biases from a network, as the autoencoders tested featured entirely isotropic primitives, except for the activation functions. This establishes isotropies’ inductive bias over a wider variety of primitives; follow-up can extend this for anisotropic networks by selectively reintroducing permutation-derived functions in other primitives.
@@
Additionally, this may suggest that interpretable models and AI safety are especially tractable due to anisotropic inductive biases. Thus, design choices between anisotropy, isotropy, quasi-isotropy Bird (2025a), or other definitions may allow tuning of models to balance these behaviours with representation capacity and performance — especially if the pathological nature is further validated. This provides a framework and implementable tunable mode, which earlier work encourages for greater interpretability and safety Elhage et al. (2022). However, it would also suggest that representations generated from isotropic implementations currently act in opposition to the assumptions of many interpretability approaches. Heuristically, this suggests that isotropic networks yield more unconstrained representations, which may be more natural or inherent to deep learning structures, as they are free from the human-imposed task-agnostic discretising structure imposed by anisotropy. Yet, it is also often this discretising structure which is made interpretable when understanding AI mechanisms. Hence, bridging such an interpretability tool gap may be beneficial, and the PPP method offers a starting point.
@@
Overall, this implicates discrete-permutation defined forms of activation functions, as stimulating a task-agnostic alignment about the standard basis. This contrasts with continuous-orthogonal defined forms, which do not. This indicates that forms carry an inductive bias for representations, which appears to be indicative of the maximal symmetry group structure of the form. Activation functions are demonstrated as a direct causal mechanism for quantised representations, and several downstream emergent phenomena are established through connection to prior work Bird (2025b). Hence, this approach has established a causal origin for this interpretable structure and several interpretability phenomena.
@@ A.2 Spontaneous Symmetry Breaking Artefacts
Often, one might expect representations to be isotropically distributed on a fresh initialisation of the model, especially if a left- and right-flavour probabilistic invariant initialiser Bird [2025a] is employed. However, spontaneous symmetry breaking upon choosing a particular initialisation can produce artefactual geometry due to a dataset’s subspace.
Since biases are initialised to zero, these initialisations produce linear, not affine, maps [^5] The success in this form of bias initialisation could be reconsidered and freshly explored under the lack of symmetry breaking perspective developed under the wider taxonomy Bird [2025a].. If a dataset is strongly anisotropic, such as being situated in a $\left[0,1\right]^{28\times 28}$ subspace, then when it is randomly embedded, it has a larger chance that specific directions are measured to have an anisotropy. For example, where the $\vec{1}$ becomes embedded will be angularly denser in representations than where any $\hat{e}_{i}$ is embedded. These embedded representational anisotropies often occur anti-aligned due to the small chance of spontaneously aligning precisely with the distinguished directions for the plot.
@@ B.1 Dependence on Network Width
Figure 8: This network shows the PPP method applied over networks of 0 hidden layer blocks, $18$ neurons and using input-output normalisation as described in Apps. A and D. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in an identical fashion to Fig. 1. Overall, this plot shows that anisotropy tends to result in the emergence of axis-aligned representation clusters, including several instances of apparent antipodal alignments. Whilst isotropic examples do not produce such a structure and tend to be more evenly distributed. It can be seen that under normalisation, there is a tendency for more diffuse clusters in anisotropic results, particularly evident in narrower networks.
@@ B.2 Dependence on Autoencoder Depth
Figure 11: This network shows the PPP method applied over networks of $1$ hidden layer block, $18$ neurons and using input-output normalisation as described in Apps. A and D. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. It is evident from this plot that the bottom-five isotropic networks tend to produce a smooth, continuous-like distribution of representations. The top-five anisotropic plots indicate the emergence of structure tending to be anti-aligned to the standard basis. This is more sparse, even than the uniform isotropy, likely since many representations are producing narrow anti-aligned beams in the anisotropic case. The PPP method is likely only capturing the more diffuse tail of such a cluster due to the $\epsilon=0.75$ thresholding, which is significantly more stringent than $2/\sqrt{18\times 2}$ required to include anti-aligned diagonals. Overall, this suggests the anisotropic networks are producing an anti-aligned structure consistent with a dense-like code.
@@ B.3 Dependence on Latent Layer of Study
Figure 14: This network shows the PPP method applied over networks of $3$ hidden layer block, $18$ neurons and using input-output normalisation as described in Apps. A and D. PPP was applied to the $0^{\text{th}}$ hidden layer. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. One can observe that representations are absent in anisotropic rows 1, 2, and 4, likely a result of anti-alignment, whereas some axis-aligned structure is present in rows 3 and 5. Isotropic rows demonstrate larger, more continuous representational structures, but also demonstrate clustering such that they are not distributed smoothly over the representation space.
@@ B.4 Results for Leaky-ReLU
In this section, the standard Leaky-ReLU, shown in Eqn. 6, is compared with its isotropic counterpart, shown in Eqn. 7 Bird [2025a].
@@ B.4.1 Hyperoctahedral Leaky-ReLU
Figure 24: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 22 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 8. Both rows 11 and $5$ indicate a tendency for representation overdensities aligning with all standard basis vectors; in the remaining rows, a pattern is not so evident, but a faint alignment remains visible, particularly in rows 22 and $3$.
@@ B.5 MNIST Results
Figure 28: Displays the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 33-hidden-layer blocks consisting of $18$ neurons per block. This architecture is comparable to those in App. D in every way except for the dimensions of CIFAR, which is exchanged for MNIST shape on input-output layers. Overall, one can observe that isotropic networks in the bottom five rows continue to produce smooth distributions that fill the representation space. In comparison, anisotropic networks on the top five rows tend to have many representations centred around the origin, and some produce axis-aligned features.

### v4

@@ 5 Conclusion
In particular, this structure appears indicative of the maximal symmetry group to which the function algebraically belongs. This is achieved through comparisons of various activation functions with differing maximal symmetries, indicating that symmetry is a good primary predictor rather than solely the specifics of the chosen function. This empirically motivates a thorough investigation of all primitives in terms of their resultant inductive biases when interacting with general networks, previously encouraged in Bird [4].
Tanh-like activation functions are chosen in two ways. The first is when their form is defined through maximally satisfying a permutation and sign-flip algebraic equivariance of the standard basis, termed the hyperoctahedral group. This function definition results in representational clusters about standard distinguished directions, which accumulate through training, progressively approximating a discrete direction. Whereas, functional forms which maximally abide by an orthogonal algebraic equivariance produce apparent smooth continua of representations. This indicates that the hyperoctahedral discrete group is inducing an approximately-discrete structure to emerge in representation space, implying the functional form is acting as a quantising bias on the representations. This also demonstrates that the prediction that isotropic functional forms eliminate task-agnostic discretising structure is validated. A similar result is found for algebraic permutation symmetry standard Leaky-ReLU, two differing hyperoctahedral and an orthogonal variation of Leaky-ReLU. These also produce a quantising bias characteristic of the discrete groups, whereas isotropic Leaky-ReLU tends to produce smooth continua. This reinforces that the algebraic symmetry is a good predictor of the inductive bias besides solely function specifics, since Tanh-like and Leaky-ReLU-like functions differ substantially in analytical qualities. This empirically validates an earlier hypothesis [4]. Additionally, these inductive biases seem broadly consistent across specific instances of groups, such as $\mathrm{O}\left(18\right)$, $\mathrm{O}\left(24\right)$, etc., indicating that the overall group family, e.g. orthogonal, can be a good categorical label for these inductive biases.
@@
Consequently, the empirical results strongly support the prediction that choices in activation function form are not benign, but impart unintended inductive biases corresponding to their symmetry properties, which are sufficient in many cases to strongly influence representations, confirming a prior prediction. The phenomenon is expected to be systematic for more general anisotropic primitive forms, and follow-up studies can investigate this.
@@
The empirical evidence thus far strongly suggests the link between algebraic symmetries and inductive biases on representations. If such a finding holds over further analyses and is determined to be generally valid over a broad scope of primitives and models, then one might conjecture that Algebraic Symmetries of Functions Utilised in Models Tend to Propagate Their Geometry Into Emergent Representational Structures Through Training as a potential principle for representational geometry. Despite the evidence presented in this paper, this general statement remains a tentative conjecture at the current time, requiring substantial further investigation. Refinement of this statement and detailed mathematical modelling can continue to be evaluated. For example, this may not be expected to hold where the symmetry cannot ‘act’ on the network, such as in the case of algebraic end-to-end symmetries [6] which use it to transform representations with the data structure — unless this is reframed as a side-effect of such an inductive bias.
@@
Additionally, combinations of network primitives are expected to exert influence on the representations of larger models, potentially in an interacting, non-trivial hierarchical manner. Understanding this is a pivotal goal of Taxonomic Deep Learning’s approach [4]. This is partially demonstrated in the deeper models discussed in the appendices. Yet, overall, these potentially conflating and obscuring influences motivated the first-principles, minimalistic, and ablation-study approach undertaken in this paper, which isolates the cause to individual primitives in otherwise discretising bias-free autoencoders, as opposed to larger models.
@@
This work also corroborates the findings from the Spotlight Resonance Method study and provides a more fundamental causal framework for its observed alignment phenomenon. Coupled with the findings of this prior study, this work suggests a function-driven, symmetry-based causal mechanism for the emergence of several interpretability phenomena, suggesting that they are predicated upon the contemporary anisotropic functional forms. This representational response to anisotropic functional forms appears to explain the downstream emergence of grandmother neurons, generally discrete linear features, axis-alignment, and the representational superposition of discrete features, amongst other phenomena. Hence, axis-alignment and grandmother neurons are shown to be phenomena stimulated by design choices, rather than being spontaneous and fundamental to deep learning as a paradigm. Whilst the other phenomena appear to be strongly dependent upon such choices. Overall, this strongly supports the emergence of axis-aligned findings that are frequently observed [3, 20] and suggests that structural asymmetry in distributions can emerge, particularly from certain functional forms. This is considered as opposing evidence for representational uniformity [25] or spontaneous production of axis-alignment. Together, these results may offer explanatory mechanisms to generalise to further representational geometry and interpretability phenomena.
@@
The isotropic deep learning paper also predicts that this phenomenon extends to semantics. These results demonstrate the formation of approximately discrete representations, which, when combined with the spotlight resonance method’s attribution of semantics to these representations, are considered to validate the semantic prediction as well. An end-to-end analysis could also be undertaken in future work. This direct mode for the formation of grandmother neurons is currently limited to deep learning and does not establish an empirical link to biological systems. Nevertheless, it could be used to formulate an interdisciplinary hypothesis.
In addition, the Privileged-Plane Projective method offers an alternative tool for determining alignment phenomena in representational spaces, while retaining both angular and radial information. This is beneficial where a visualisation of the complicated geometry of representation distributions is desirable. The prior SRM may still be better suited to statistical analysis techniques, although it comes at the cost of reduced angular fidelity and no radial information. If suitably discerning quantitative statistical metrics can be developed, they may be an advantageous future direction for supporting the conclusions drawn. A further promising direction may be to construct a hybrid methodology that incorporates the benefits and mitigates the shortcomings of each method in a joint, visualisable, and statistical tool. Additionally, once the induced geometry is better understood, one could derive suitable distribution estimates for anisotropic and isotropic representations, enabling statistical measures to yield further insight into the geometry. At this time, predicting such distributions is not evident from the theory, as the specific positioning of clusters depends on the function in use. However, their overall organisation is functional-form induced. This would also require a quantity to determine the angular ‘usefulness’ of the function. This may reinforce conclusions, but will need follow-up work to assess the appropriate implementation of such a joint approach. The visualisations are still considered sufficient for the conclusions drawn, given the clear, distinctive, and repeated occurrences of the characteristic symmetry-organised signatures across a range of setups. However, statistical measures may further strengthen the certainty of such an interpretation if an appropriately rigorous methodology can be developed to discern these signatures numerically.
@@
The presence of discrete-like structure is also suggested to be often circumstantially pathological [4]. Preliminary results support this statement for autoencoding networks. Isotropic formulations are shown to enable the systematic elimination of these biases from a network, as the autoencoders tested featured entirely isotropic primitives, except for the activation functions. This establishes isotropies’ inductive bias over a wider variety of primitives; follow-up can extend this for anisotropic networks by selectively reintroducing permutation-derived functions in other primitives.
@@
Additionally, this may suggest that interpretable models and AI safety are especially tractable due to anisotropic inductive biases. Thus, design choices between anisotropy, isotropy, quasi-isotropy [4], or other definitions may allow tuning of models to balance these behaviours with representation capacity and performance — especially if the pathological nature is further validated. This provides a framework and implementable tunable mode, which earlier work encourages for greater interpretability and safety [8]. However, it would also suggest that representations generated from isotropic implementations currently act in opposition to the assumptions of many interpretability approaches. Heuristically, this suggests that isotropic networks yield more unconstrained representations, which may be more natural or inherent to deep learning structures, as they are free from the human-imposed task-agnostic discretising structure imposed by anisotropy. Yet, it is also often this discretising structure which is made interpretable when understanding AI mechanisms. Hence, bridging such an interpretability tool gap may be beneficial, and the PPP method offers a starting point.
@@
Overall, this implicates discrete-permutation defined forms of activation functions, as stimulating a task-agnostic alignment about the standard basis. This contrasts with continuous-orthogonal defined forms, which do not. This indicates that forms carry an inductive bias for representations, which appears to be indicative of the maximal symmetry group structure of the form. Activation functions are demonstrated as a direct causal mechanism for quantised representations, and several downstream emergent phenomena are established through connection to prior work [5]. Hence, this approach has established a causal origin for this interpretable structure and several interpretability phenomena.
@@ A.2 Spontaneous Symmetry Breaking Artefacts
Often, one might expect representations to be isotropically distributed on a fresh initialisation of the model, especially if a left- and right-flavour probabilistic invariant initialiser [4] is employed. However, spontaneous symmetry breaking upon choosing a particular initialisation can produce artefactual geometry due to a dataset’s subspace.
Since biases are initialised to zero, these initialisations produce linear, not affine, maps [^9] The success in this form of bias initialisation could be reconsidered and freshly explored under the lack of symmetry breaking perspective developed under the wider taxonomy [4].. If a dataset is strongly anisotropic, such as being situated in a $\left[0,1\right]^{28\times 28}$ subspace, then when it is randomly embedded, it has a larger chance that specific directions are measured to have an anisotropy. For example, where the $\vec{1}$ becomes embedded will be angularly denser in representations than where any $\hat{e}_{i}$ is embedded. These embedded representational anisotropies often occur anti-aligned due to the small chance of spontaneously aligning precisely with the distinguished directions for the plot.
@@ B.1 Dependence on Network Width
Figure 8: This network shows the PPP method applied over networks of 0 hidden layer blocks, $18$ neurons and using input-output normalisation as described in Apps. A and E. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in an identical fashion to Fig. 1. Overall, this plot shows that anisotropy tends to result in the emergence of axis-aligned representation clusters, including several instances of apparent antipodal alignments. Whilst isotropic examples do not produce such a structure and tend to be more evenly distributed. It can be seen that under normalisation, there is a tendency for more diffuse clusters in anisotropic results, particularly evident in narrower networks.
@@ B.2 Dependence on Autoencoder Depth
Figure 11: This network shows the PPP method applied over networks of $1$ hidden layer block, $18$ neurons and using input-output normalisation as described in Apps. A and E. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. It is evident from this plot that the bottom-five isotropic networks tend to produce a smooth, continuous-like distribution of representations. The top-five anisotropic plots indicate the emergence of structure tending to be anti-aligned to the standard basis. This is more sparse, even than the uniform isotropy, likely since many representations are producing narrow anti-aligned beams in the anisotropic case. The PPP method is likely only capturing the more diffuse tail of such a cluster due to the $\epsilon=0.75$ thresholding, which is significantly more stringent than $2/\sqrt{18\times 2}$ required to include anti-aligned diagonals. Overall, this suggests the anisotropic networks are producing an anti-aligned structure consistent with a dense-like code.
@@ B.3 Dependence on Latent Layer of Study
Figure 14: This network shows the PPP method applied over networks of $3$ hidden layer block, $18$ neurons and using input-output normalisation as described in Apps. A and E. PPP was applied to the $0^{\text{th}}$ hidden layer. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. One can observe that representations are absent in anisotropic rows 1, 2, and 4, likely a result of anti-alignment, whereas some axis-aligned structure is present in rows 3 and 5. Isotropic rows demonstrate larger, more continuous representational structures, but also demonstrate clustering such that they are not distributed smoothly over the representation space.
@@ B.4 Results for Leaky-ReLU
In this section, the standard Leaky-ReLU, shown in Eqn. 6, is compared with its isotropic counterpart, shown in Eqn. 7 [4].
@@ B.4.1 Hyperoctahedral Leaky-ReLU
Figure 24: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 22 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 8. Both rows 11 and 55 indicate a tendency for representation overdensities aligning with all standard basis vectors; in the remaining rows, a pattern is not so evident, but a faint alignment remains visible, particularly in rows 22 and $3$.
@@ B.5 MNIST Results
Figure 28: Displays the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 33-hidden-layer blocks consisting of $18$ neurons per block. This architecture is comparable to those in App. E in every way except for the dimensions of CIFAR, which is exchanged for MNIST shape on input-output layers. Overall, one can observe that isotropic networks in the bottom five rows continue to produce smooth distributions that fill the representation space. In comparison, anisotropic networks on the top five rows tend to have many representations centred around the origin, and some produce axis-aligned features.